40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13280     Entry time: Thu Aug 31 00:52:52 2017     In reply to: 13276     Reply to this: 13281
Author: gautam 
Type: Update 
Category: LSC 
Subject: REFL55 whitening board debugging 

[rana,gautam]

We did an ingenious checkup of the whitening board tonight.

  • The board is D990694
  • We made use of a tip-tilt DAC channel for this test (specifically TT1 UL, which is channel 1 on the AI board). We disconnected the cable going from the AI board to the TT coil driver board.
    • as opposed to using a function generator to drive the whitening filter, this approach allows us to not have to worry the changing offsets as we switch the whitening gain.
    • By using the CDS system to generate the signal and also demodulate it, we also don't have to worry about the drive and demod frequencies falling out of sync with each other.
  • The test was done by injecting a low frequency (75.13 Hz, amplitude=0.1) excitation to this DAC channel, and using the LSC sensing matrix infrastructure to demodulate REFL55 I and Q at this frequency. Demod phases in these servos were adjusted such that the Q phase demodulated signal was minimized.
  • An excitation was injected using awggui into TT1 UL exc channel.
  • We then stepped the whitening gains for REFL55_I and REFL55_Q in 3dB steps, waiting 5 seconds for each step. Syntax is z step -s 5 C1:LSC-REFL55_I_WhiteGain +1.0,15 C1:LSC-REFL55_Q_WhiteGain +1.0,15
  • Attachment #1 suggests that the whitening filter board is working as expected (each step is indeed 3dB and all steps are equal to the eye).
  • Data + script used to generate this plot is in Attachment #2.

I've restored all connections at that we messed with at the LSC rack to their original positions.

The TT alignment seems to be drifting around more than usual after we disconnected one of the channels - when I came in today afternoon, the spot on the AS camera had drifted by ~1 spot diameter so I had to manually re-align TT1. 

Quote:
 

Based on my tests, everything on the Demod board seems to work as expected. I need to think more about what else could be happening here - specifically do a more direct test on the whitening board.

Attachment 1: REFL55_whtCheck.pdf  25 kB  Uploaded Thu Aug 31 02:13:33 2017  | Hide | Hide all
REFL55_whtCheck.pdf
Attachment 2: REFL55_whtChk.tar.gz  21 kB  Uploaded Thu Aug 31 02:15:39 2017
ELOG V3.1.3-